# Pretraining Language Models to Ponder in Continuous Space

### v2

@@ 3.1 Small Scale Validation on GPT-2 and LLaMA
The detailed configurations of the models, including sizes, learning rates, and batch sizes, are specified in Table 1. These hyperparameters primarily follow the GPT-3 specifications (Brown et al., 2020). However, unlike GPT-3, we untie the input and output embedding matrices.
@@
Figure 2: (top) Scaling curves of GPT3 LLaMA and their corresponding pondering models. (bottom) Relative improvements of RoPE + RMSNorm + SwiGLU MLP and Pondering.
Results. Figure 2 (top) illustrates the scaling curves of the validation loss on the Pile subset for vanilla GPT-2, vanilla LLaMA, and their pondering counterparts. Our results show that incorporating pondering significantly improves the performance for both GPT-2 and LLaMA across the entire size range tested (405M to 1.4B parameters). For instance, by fitting scaling curves, we find that PonderingGPT-834M achieves a loss comparable to a vanilla GPT-2 model trained with approximately 2.01×2.01$\times$2.01 × parameters*tokens, while PonderingLLaMA-834M matches the loss of vanilla LLaMA trained with roughly 2.26×2.26$\times$2.26 × parameters*tokens.
Furthermore, Figure 2 (bottom) shows the relative validation loss improvements (denoted as $\Delta$ loss) of architectural modifications inherent in LLaMA—namely rotary positional embeddings (RoPE) (Su et al., 2024), RMSNorm (Zhang & Sennrich, 2019), and SwiGLU activation—in comparison to GPT-2, alongside the relative improvements obtained by our pondering method for both architectures. We observe that while the relative improvement due to RoPE, RMSNorm, and SwiGLU MLP diminishes as model size and data scale increase (reaching approximately 0.02 for the 1.4B parameter model), the relative improvement provided by pondering consistently remains around 0.1 across the entire tested scale.
Figure 3: Language modeling loss when scaling parameter count and training tokens. PonderingPythia achieves comparable performance to the official Pythia-6.9B while using only 37% of the parameters, and matches the performance of the official Pythia-1B with just 41% of the training tokens.
Table 2: Language Modeling Perplexity (ppl). Lower is better. The values in parentheses indicate the improvement ($\downarrow$) compared to the corresponding Pythia model. For simplicity, we refer to PonderingPythia as Ponder.
Model Pile Wikitext Lambada Openai Lambada Standard
Pythia-70M $18.36$ $57.03$ $141.94$ $967.72$
Ponder-70M $12.68$($\downarrow$ 5.68) $34.07$($\downarrow$ 22.96) $39.19$($\downarrow$ 102.75) $145.51$($\downarrow$ 822.21)
Pythia-160M $12.55$ $33.43$ $38.15$ $186.51$
Ponder-160M $9.87$($\downarrow$ 2.68) $23.69$($\downarrow$ 9.74) $15.96$($\downarrow$ 22.19) $41.24$($\downarrow$ 145.27)
Pythia-410M $8.85$ $20.11$ $10.86$ $31.54$
Ponder-410M $8.15$($\downarrow$ 0.70) $17.78$($\downarrow$ 2.33) $8.55$($\downarrow$ 2.31) $17.48$($\downarrow$ 14.06)
Pythia-1B $7.85$ $16.45$ $7.91$ $17.41$
Ponder-1B $7.15$($\downarrow$ 0.70) $14.61$($\downarrow$ 1.84) $6.02$($\downarrow$ 1.89) $10.64$($\downarrow$ 6.77)
Pythia-1.4B $7.24$ $14.72$ $6.08$ $10.87$
Ponder-1.4B $6.82$($\downarrow$ 0.42) $13.59$($\downarrow$ 1.13) $5.37$($\downarrow$ 0.71) $8.88$($\downarrow$ 1.99)
Pythia-2.8B $6.63$ $12.69$ $5.04$ $8.23$
Ponder-2.8B $6.21$($\downarrow$ 0.42) $11.74$($\downarrow$ 0.95) $4.15$($\downarrow$ 0.89) $6.17$($\downarrow$ 2.06)
Pythia-6.9B $6.29$ $11.41$ $4.45$ $6.95$
3.2 Large-Scale Pretraining on Pile
We further validate the effectiveness of our pondering method by conducting large-scale pretraining experiments on the entire Pile dataset (300B tokens) (Gao et al., 2020). We train a new model, named PonderingPythia, from scratch using exactly the same architectural components (parallel attention and MLP layers, rotary embeddings with 1/4 head dimensions), same tokenizer and training hyperparameters (optimizer settings, learning rate schedule, batch size, and context length) as the original Pythia models. We then compare PonderingPythia models’ scaling curves and language modeling capabilities with the official Pythia model suite.
3.2.1 Scaling Curves
We plot the PonderingPythia and official Pythia model’s scaling curves along parameters size, training tokens and training flops. As depicted in Figure 3 (left), the fitted curves show that a 2.5B-parameter PonderingPythia model achieves a validation loss comparable to the 6.9B-parameter official Pythia model, while requiring only 37% of the parameters. In Figure 3 (middle), we evaluated the 1B PonderingPythia and official Pythia models at intervals of 20B training tokens. The fitted curves indicate that PonderingPythia trained with 115B tokens achieves performance comparable to the official Pythia model trained with 280B tokens, consuming only 41% of the training tokens. In Figure 3 (right), we report the language modeling loss of vanilla Pythia-70M [^4] To match the training FLOPs of vanilla Pythia-70M with PonderingPythia-70M, we trained the vanilla model for 4 epochs. and PonderingPythia-70M under the same computational budget during pretraining. It can be observed that PonderingPythia-70M consistently outperforms vanilla Pythia-70M when trained with the same number of FLOPs.
3.2.2 Language Modeling Ability Evaluation
We measure the perplexity on several language modeling datasets to reflect general language modeling capabilities. Specifically, we report perplexity scores on the Pile validation set, Wikitext, and the Lambada dataset (both OpenAI and standard versions), detailed in Table 2. The results demonstrate significant perplexity improvements across all datasets and model sizes. Notably, the perplexity achieved by the PonderingPythia-2.8B model is even better than that of the official Pythia-6.8B model, which is nearly 2.5 times larger.
3.3 Downstream Tasks Evaluation
We use the PonderingPythia models pretrained in the previous subsection to conduct downstream task evaluations.
Figure 4: Instruction-following abilities evaluated on MT-Bench. PonderingPythia-1B and 1.4B consistently outperform their corresponding official Pythia models across all subtasks.
3.3.1 General Downstream Tasks
We consider various widely-used benchmarks, including the tasks originally used by Pythia (LAMBADA (Paperno et al., 2016), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), SciQ (Welbl et al., 2017). Additionally, we include HellaSwag (Zellers et al., 2019) for commonsense reasoning and RACE (Lai et al., 2017) for reading comprehension.
We evaluate both zero-shot and five-shot learning performance using the LM evaluation harness (Gao et al., 2023). Detailed results are shown in Table 3. Across all evaluated model sizes, PonderingPythia consistently and significantly outperforms the official Pythia models, as well as comparable OPT, GPT-Neo, and Bloom models. Remarkably, with only 1/10 of the training data (300B tokens) and fewer parameters (1B vs. 1.1B), our PonderingPythia-1B achieves results comparable to, or even surpassing, TinyLlama—which uses a more advanced LLaMA architecture and 3T tokens. PonderingPythia-2.8B also surpasses Pythia-6.9B, which is nearly 2.5 times its size.
Table 3: Five-shot and zero-shot evaluations on downstream NLP tasks. All pretrained model weights used for comparison are obtained from their official repositories. We refer to PonderingPythia as Ponder for simplicity. $\Delta$acc is compared to the official Pythia models.
Model (#training tokens) Lambada OpenAI ARC -E Lambada Standard ARC -C Wino Grande PIQA Hella Swag SciQ RACE Avg acc / $\Delta$acc $\uparrow$
@@ 3.3.1 General Downstream Tasks
Pythia-70M (300B) 11.9 36.7 9.2 17.1 50.5 58.7 26.7 57.8 25.1 32.6
Ponder-70M (300B) 28.4 44.6 21.1 18.6 52.2 62.4 28.4 75.7 26.9 39.8 / +7.2
Pythia-160M (300B) 24.9 44.7 19.0 18.4 50.4 63.5 28.6 76.4 27.8 39.3
OPT-125M (300B) 29.8 42.3 26.4 19.7 51.3 63.4 29.0 78.8 30.1 41.2
GPTneo-125M (300B) 29.3 43.9 23.0 19.1 51.9 63.9 28.5 79.9 27.9 40.8
Ponder-160M (300B) 39.9 50.1 31.2 22.1 51.2 65.8 31.9 85.4 30.1 45.3 / +6.0
Pythia-410M (300B) 43.9 54.7 32.8 22.3 53.4 68.0 33.8 88.9 30.4 47.6
OPT-350M (300B) 38.3 45.4 32.1 20.5 53.0 65.8 31.9 85.7 29.5 44.7
Bloom-560M (366B) 29.4 50.2 29.7 21.9 52.7 64.2 31.4 88.0 30.0 44.2
Ponder-410M (300B) 48.9 58.7 43.7 26.1 54.0 70.5 37.3 91.0 32.4 51.4 /+3.8
Pythia-1B (300B) 48.3 58.6 35.8 25.4 52.8 71.3 37.7 91.6 31.7 50.4
OPT-1.3B (300B) 54.0 60.4 49.0 26.9 56.9 72.4 38.5 91.8 35.4 52.7
GPTneo-1.3B (300B) 49.9 59.9 44.5 25.9 56.6 71.6 38.6 86.1 34.5 51.9
Bloom-1.1B (366B) 36.3 54.9 37.4 24.9 53.4 67.6 34.8 88.7 33.0 47.9
Ponder-1B (300B) 57.7 63.2 52.5 28.6 58.6 73.3 41.9 93.4 36.3 56.2 / +5.8
Pythia-1.4B (300B) 54.5 63.1 44.5 28.8 57.1 71.0 40.5 92.4 34.6 54.1
Bloom-1.7B (366B) 42.5 58.8 41.5 26.2 57.7 68.7 37.6 91.9 33.5 50.9
Ponder-1.4B (300B) 59.2 67.5 49.9 32.4 60.4 73.5 44.2 94.3 37.1 57.6 / +3.5
Tinyllama-1.1B (3T) 53.8 64.8 45.0 31.1 59.4 73.8 44.9 94.0 36.4 55.9
OPT-2.7B (300B) 60.2 64.7 55.0 29.8 62.2 75.1 46.1 93.0 37.5 58.2
GPTneo-2.7B (300B) 56.0 64.0 51.6 30.1 59.6 73.9 42.4 93.3 35.5 56.3
Bloom-3B (366B) 46.2 63.8 47.1 31.7 57.8 70.8 41.4 93.4 34.6 54.1
Pythia-2.8B (300B) 59.0 67.0 50.7 31.0 61.1 74.4 45.3 93.7 35.9 57.6
Pythia-6.9B (300B) 62.5 69.6 54.8 35.6 62.9 76.6 48.0 94.6 36.7 60.1
Ponder-2.8B (300B) 64.2 70.6 58.7 35.8 65.3 76.7 49.0 94.3 39.0 61.5 / +3.9
@@
Pythia-70M (300B) 18.7 36.7 13.5 18.8 51.1 59.9 26.6 59.9 23.9 34.3
Ponder-70M (300B) 33.5 42.5 24.1 19.1 51.1 61.5 28.3 70.3 27.8 39.8 / +5.5
Pythia-160M (300B) 33.0 43.8 21.4 18.9 52.0 62.2 28.4 73.7 27.9 40.1
OPT-125M (300B) 37.9 43.5 29.0 19.0 50.3 62.8 29.2 75.2 30.1 41.9
GPTneo-125M (300B) 37.5 43.9 26.0 19.1 50.4 63.0 28.7 76.5 27.5 41.4
Ponder-160M (300B) 44.3 46.8 33.8 20.5 52.3 63.9 31.9 76.8 29.6 44.4 / +4.3
Pythia-410M (300B) 51.4 52.2 36.4 21.4 53.8 66.9 33.7 81.5 30.9 47.6
OPT-350M (300B) 45.2 44.0 35.8 20.7 52.3 64.5 32.0 74.9 29.8 44.4
Bloom-560M (366B) 34.3 47.5 33.3 22.4 51.5 63.8 31.5 80.3 30.5 43.9
Ponder-410M (300B) 56.9 51.9 45.3 22.6 56.0 68.7 37.0 81.4 33.8 50.4 / +2.8
Pythia-1B (300B) 55.9 56.8 42.0 24.2 52.5 70.5 37.7 83.3 32.7 50.6
OPT-1.3B (300B) 57.9 57.1 52.5 23.4 59.7 71.8 41.6 84.3 34.3 53.6
GPTneo-1.3B (300B) 57.1 56.2 45.3 23.2 55.0 71.2 38.6 86.1 34.5 51.9
Bloom-1.1B (366B) 42.6 51.5 42.9 23.6 54.9 67.3 34.5 83.6 32.6 48.2
Ponder-1B (300B) 62.3 60.5 51.9 27.0 56.5 72.2 41.8 87.4 35.4 55.0 / +4.4
Pythia-1.4B (300B) 61.6 60.4 49.7 25.9 57.5 70.8 40.4 86.4 34.1 54.1
Bloom-1.7B (366B) 46.2 56.4 44.5 23.7 56.8 68.5 37.5 85.0 33.2 50.2
Ponder-1.4B (300B) 65.2 62.0 53.8 27.0 60.1 72.6 44.0 89.0 35.2 56.5 / +2.4
Tinyllama-1.1B (3T) 58.8 60.3 49.3 28.0 59.0 73.3 45.0 88.9 36.4 55.4
OPT-2.7B (300B) 63.5 60.8 56.0 26.8 61.2 73.8 45.9 85.8 36.2 56.7
GPTneo-2.7B (300B) 62.1 61.2 51.6 27.5 57.8 72.3 42.7 89.3 35.1 55.5
Bloom-3B (366B) 51.7 59.4 50.9 28.0 58.7 70.8 41.4 88.8 35.2 53.9
Pythia-2.8B (300B) 64.6 64.4 54.3 29.5 60.2 73.8 45.4 88.5 34.9 57.3
Pythia-6.9B (300B) 67.2 67.3 55.9 31.4 61.0 75.2 48.1 89.3 36.9 59.1
Ponder-2.8B (300B) 68.9 66.5 60.8 32.5 63.6 75.0 48.6 91.0 36.5 60.4 / +3.1
3.3.2 Instruction-following Ability Evaluation
To assess the instruction-following capability of our model, we further fine-tuned PonderingPythia-1B and 1.4B, as well as the corresponding official Pythia models, on the Alpaca dataset using the same settings (Taori et al., 2023). The fine-tuned models were evaluated with MT-Bench (Zheng et al., 2023), a popular multi-turn question benchmark. The experimental results are shown in Figure 4. As illustrated, both PonderingPythia-1B and 1.4B consistently outperform their official Pythia counterparts across all subtasks, achieving average improvements of 0.33 and 0.55, respectively. [^5] The marginal gains on Coding and Math tasks may be attributed to the limited coding and mathematical abilities of the Pythia models.
Figure 5: Increasing the number of pondering steps consistently reduces loss on the Pile validation set.
3.4 Impact of Pondering Steps
To further investigate the effect of pondering steps on model performance, we trained several 160M-parameter Pythia models from scratch with different numbers of pondering steps: 0 (baseline), 1, 3, 5, and 10. These models were pretrained on a 30B-token subset of the Pile dataset. The results, presented in Figure 5, clearly show that increasing the number of pondering steps consistently reduces the language modeling loss on the Pile validation set. This demonstrates the significant potential and scalability of our method, suggesting that model performance can be further improved by increasing the depth of pondering steps.
4 Limitations and Future Work
4.1 Limitations
4.2 Future Work
There are several promising directions for future work. Firstly, our proposed method is not limited to decoder-only architectures or language modeling; it has the potential to be applied to a wide range of model types and domains. For example, it could be adapted to state-space models such as Mamba (Gu & Dao, 2023), encoder models, or RWKV (Peng et al., 2023), as well as extended to other areas. Another promising direction is the introduction of token-adaptive pondering, which may significantly reduce computation and further enhance our method. It would also be interesting to investigate the interpretability of the pondering process, such as how the model "thinks" during pondering, the semantics of the pondering embedding, and whether the model learns to reflect on its predictions through pondering. Finally, exploring the combination of our method with orthogonal approaches such as CoT and other test-time scaling methods could also be an interesting direction.
5 Related Work
5.1 Test-Time Compute Scaling
Scaling test-time computation has emerged as an influential approach, providing an effective alternative to merely increasing model parameters (Snell et al., 2024). Current methods for enhancing test-time computation primarily fall into two categories: parallel scaling and sequential scaling (Zeng et al., 2024; Muennighoff et al., 2025).
Parallel scaling methods involve generating multiple candidate solutions simultaneously and selecting the best solution based on certain evaluation criteria. Prominent examples of this paradigm include Best-of-N search (Cobbe et al., 2021; Sun et al., 2024; Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024) and Majority Vote (Wang et al., 2022). The primary difference among these parallel approaches lies in their strategy for selecting the optimal outcome. Nevertheless, parallel scaling approaches often encounter difficulty accurately identifying the best candidate solution among multiple generated possibilities (Stroebl et al., 2024; Hassid et al., 2024).
Sequential scaling methods, conversely, focus on progressively refining the model’s reasoning through multiple iterative steps. Representative methods include CoT (Wei et al., 2022; Nye et al., 2021), iterative rethinking and revision (Huang et al., 2022; Min et al., 2024; Madaan et al., 2024; Wang et al., 2024b, b; Lee et al., 2025; Hou et al., 2025; Muennighoff et al., 2025; Li et al., 2025). Following the emergence of powerful OpenAI o1 and DeepSeek R1, leveraging extensive chain-of-thought prompting to scale test-time computation has become increasingly popular. Nonetheless, existing methods often come with limitations, including reliance on specially curated datasets (Allen-Zhu & Li, 2023), extended context windows (Zhu et al., 2025), and complex reinforcement learning strategies (Pang et al., 2025). In contrast, our method effectively circumvents these limitations.
5.2 Latent Thinking in Language Models
Latent thinking in LMs typically refers to the hidden computational processes within transformer architectures (Yang et al., 2024; Biran et al., 2024). Prior works exploring latent thinking can generally be divided into two categories based on their implementation methods:
Adding Additional Tokens: Wang et al. (2024a) proposed predicting a planning token as a discrete latent variable prior to generating reasoning steps. Similarly, Pfau et al. (2024) investigated filler tokens (e.g., “…”) and found them effective for parallelizable problems, although they also highlighted limitations compared to CoT methods, potentially constraining scalability to more complex reasoning tasks. Zhou et al. (2024) pretrained models by randomly inserting a learnable token into the training corpus, demonstrating improvements across various tasks. Furthermore, Quiet-STaR employs reinforcement learning to train models to generate intermediate rationales at each token, thereby enhancing reasoning capabilities (Zelikman et al., 2024).
Reusing Hidden States: Early Universal Transformers (Csordás et al., ) reused hidden states through a transition network, aiming to enhance the encoder-decoder transformer architecture toward Turing completeness. A more recent method(Geiping et al., 2025) iterates over multiple layers of a language model to refine intermediate hidden states, treating this process as a form of latent thinking. Giannou et al. (2023) proposed looped transformer, recycling output hidden states back into input embeddings for algorithmic tasks. Similarly, Hao et al. (2024) fine-tuned models to reason directly within continuous latent spaces, using final hidden states as embeddings to achieve reasoning without explicit CoT. In contrast, our method relies on pondering embeddings derived from the predicted probabilities of LMs.
6 Conclusion
In this paper, we introduce the pondering process into language models through solely self-supervised learning. Pondering LM can be naturally learned through pretraining on large-scale general corpora. Our extensive experiments across three widely adopted architectures—GPT-2, Pythia, and LLaMA—highlight the effectiveness and generality of our proposed method. Notably, our PonderingPythia consistently outperforms the official Pythia model on language modeling tasks, scaling curves, downstream tasks, and instruction-following abilities when pretrained on the large-scale Pile dataset. As increasing the number of pondering steps further improves language model performance, we posit that our approach introduces a promising new dimension along which language model capabilities can be scaled.

### v3

@@ Appendix B Training Details of the Scaling Law
In Section 3.1, we have discussed the basic scaling law of our pondering models. These hyperparameters primarily follow the GPT-3 specifications (brown2020language). However, unlike GPT-3, we untie the input and output embedding matrices. We specify the parameters of our models and the training hyperparameters in Table 4.
@@
Appendix C Limitations and Future Work
C.1 Limitations
C.2 Future Work
There are several promising directions for future work. Firstly, our proposed method is not limited to decoder-only architectures or language modeling; it has the potential to be applied to a wide range of model types and domains. For example, it could be adapted to state-space models such as Mamba (gu2023mamba), encoder models, or RWKV (peng2023rwkv), as well as extended to other areas. Another promising direction is the introduction of token-adaptive pondering, which may significantly reduce computation and further enhance our method. It would also be interesting to investigate the interpretability of the pondering process, such as how the model "thinks" during pondering, the semantics of the pondering embedding, and whether the model learns to reflect on its predictions through pondering. Finally, exploring the combination of our method with orthogonal approaches such as CoT and other test-time scaling methods could also be an interesting direction.
Figure 8: Training loss during continual pretraining. PonderPythia-1B achieves a lower loss than the vanilla baseline, suggesting greater learning efficiency.
Appendix D Integrating Pondering via Continual Pretraining
To investigate if our pondering mechanism can be integrated into existing large language models, we conducted a continual pretraining experiment. We started with the pre-trained Pythia-1B model and continued its training on a 30-billion-token subset of The Pile. We compare two approaches: a standard continual pretraining of the vanilla model (Pythia-1B-Vanilla-CPT) and continual pretraining with our pondering mechanism (PonderPythia-1B-CPT).
Our results show that this integration is effective. As seen in the training loss curves in Figure 8, PonderPythia consistently achieves a lower loss, demonstrating superior learning efficiency. This efficiency translates to improved downstream performance, as detailed in Table 5. The PonderPythia model achieves a higher average accuracy in both 5-shot and 0-shot settings across a suite of benchmarks.
Collectively, these findings confirm that continual pretraining is a viable and effective strategy for equipping pre-trained models with our pondering mechanism, enhancing both adaptation efficiency and downstream capabilities.
Table 5: Downstream task performance after continual pretraining on 30B tokens. PonderPythia-1B shows improved average accuracy in both 5-shot and 0-shot evaluations.
[…]
@@ Appendix D Integrating Pondering via Continual Pretraining
Pythia-1B-Vanilla-CPT 49.4 60.0 36.5 27.1 51.9 71.4 37.9 90.4 31.9 50.7
PonderPythia-1B-CPT 49.5 58.8 42.6 25.4 54.3 69.2 37.9 91.2 34.7 51.5 (+0.8)
@@
Pythia-1B-Vanilla-CPT 56.7 56.3 42.7 24.6 52.5 70.6 37.7 83.6 32.6 50.8
PonderPythia-1B-CPT 56.7 56.3 45.4 23.8 55.6 68.3 37.8 86.3 33.0 51.5(+0.7)
Appendix E Scaling Laws with Different Pondering Steps
To further validate the robustness of our method, we extended our analysis to LLaMA models trained with different numbers of pondering steps. Specifically, we tested configurations with 2 and 4 steps while keeping all other experimental settings constant. As illustrated in Figure 9, our method remains consistently effective, with model performance improving further as the number of pondering steps increases.
Figure 9: Scaling curves of LLaMA with different pondering steps.
Appendix F Analysis on Input Embeddings and Output Distributions
To better understand the dynamics of the pondering process, we analyze the properties of input embeddings and output distributions using the Pythia-70M model trained with 10 pondering steps (as described in Section 4.2). We conduct the evaluation on a batch of 64 validation sequences with a context length of 2048 tokens.
F.1 Input Embeddings
We first examine the evolution of the input embeddings $E^{s}$ across pondering steps.
Cosine Similarity.
To measure the convergence of the embeddings, we calculate the cosine similarity between consecutive steps, $Cosine(E^{s-1},E^{s})$. Here we define $s=0$ as the original input embedding (before pondering). As shown in Figure 11, the similarity starts at approximately 0.88 and gradually converges to $1.0$, indicating that the input representations stabilize over time.
Figure 10: The cosine similarity between consecutive embedding states $E^{s-1}$ and $E^{s}$ across pondering steps.
Figure 11: KL divergence between consecutive output distributions, averaged over tokens.
Spectral Properties.
We further analyze the geometry of the embeddings. Figure 12 visualizes the spectral energy distribution (explained variance ratio) of the top singular values. Consistent with this distribution, we observe a steady decrease in Effective Rank and a simultaneous increase in Cumulative Variance (Figure 12). This collective evidence indicates that the embedding energy progressively concentrates into dominant components during the pondering process.
Figure 12: Left: Spectral energy (explained variance ratio, log scale) of the top-8 singular values of the token-embedding matrix at each pondering step $s$. Right: Effective rank ($\downarrow$ more spiky) and cumulative variance of the top-8 components ($\uparrow$ more spiky) versus step $s$, indicating progressive concentration of energy into few dominant directions.
F.2 Output Distributions
We also monitor the changes in the model’s predictions by calculating the Kullback-Leibler (KL) divergence between the output probability distributions $P^{s}$ of consecutive steps, $D_{KL}(P^{s-1}\|P^{s})$.
As illustrated in Figure 11, the KL divergence exhibits large variations in the initial steps (starting around 0.35) and decreases monotonically thereafter, suggesting that the model’s predictions refine quickly and then settle into a stable state.
Appendix G Gradient Norms Comparison
In this section, we take a closer look at training stability by comparing the gradient norms of the PonderLM against the vanilla GPT-2 1.4B model (Section 3.1). Both models were trained from scratch on the Pile using the exact same data and hyperparameters. As shown in Figure 13, while the pondering model shows a few minor spikes early on, it quickly recovers and maintains a stable training trajectory similar to the baseline.
Figure 13: Gradient norms during pre-training for vanilla vs. PonderLM 1.4B.
Appendix H Case Studies
To provide a deeper insight into the internal pondering process of PonderLM, we present case studies derived from the pre-trained PonderPythia-2.8B model. The following tables visualize the evolution of the top-3 candidate tokens and their corresponding probabilities across sequential pondering steps.
The speed of light in vacuum is approximately 3 times ten to the power of
Output Probs Pondering steps 1 Pondering steps 2 Pondering steps 3 Final predicted
Rank 1 logarithm (0.45) ˆ- (0.38) _8 (0.48) 8 (0.55)
Rank 2 , (0.38) aggreg (0.32) 8 (0.36) eight (0.23)
Rank 3 ten (0.18) approximately (0.31) tion (0.17) 10 (0.22)
The chemical symbol for silver is
Output Probs Pondering steps 1 Pondering steps 2 Pondering steps 3 Final predicted
Rank 1 symbol (0.47) atoms (0.39) Ag (0.40) Ag (0.94)
Rank 2 symbols (0.32) atomic (0.32) symbol (0.33) " (0.04)
Rank 3 nickname (0.20) elemental (0.30) elements (0.27) S (0.02)
The largest ocean on Earth is the
Output Probs Pondering steps 1 Pondering steps 2 Pondering steps 3 Final predicted
Rank 1 oceans (0.66) Pacific (0.40) Pacific (0.59) Pacific (0.77)
Rank 2 ocean (0.32) oceans (0.32) oceans (0.21) Atlantic (0.19)
Rank 3 seas (0.02) Antarctic (0.28) ocean (0.20) Indian (0.03)
The derivative of sin x is
Output Probs Pondering steps 1 Pondering steps 2 Pondering steps 3 Final predicted
Rank 1 differentiable (0.55) homework (0.61) approximately (0.37) cos (0.36)
Rank 2 the (0.23) differentiable (0.21) cos (0.32) x (0.34)
Rank 3 derivatives (0.21) basics (0.17) differentiable (0.31) (0.30)
The opposite of north is
Output Probs Pondering steps 1 Pondering steps 2 Pondering steps 3 Final predicted
Rank 1 directions (0.41) directions (0.40) south (0.92) south (0.95)
Rank 2 compass (0.36) noun (0.32) unclear (0.04) east (0.02)
Rank 3 the (0.22) idiot (0.28) South (0.03) not (0.02)
The chemical symbol for gold is
Output Probs Pondering steps 1 Pondering steps 2 Pondering steps 3 Final predicted
Rank 1 symbol (0.48) atoms (0.45) _Au (0.78) Au (0.87)
Rank 2 symbols (0.40) elemental (0.28) Au (0.17) " (0.09)
Rank 3 approximately (0.12) metals (0.27) element (0.05) the (0.04)